# Adaptive Learn-then-Test: Statistically Valid and Efficient Hyperparameter Selection

### v1

@@ 5.2.1 Problem Definition
[…]
@@ 5.2.2 Results
We study a setting with differentiated service requirements. In particular, we enforce a constraint on the average queue occupancy and energy efficiency for high-priority users, while making a best-effort attempt at minimizing the energy-delay product for low-priority users. Writing as $Q^{\rm HI}(\lambda)=\mathbb{E}_{P_{Z}}[Q^{\rm HI}(\lambda,Z)]$ the average queue occupancy of high-priority users, and as $\rho^{\rm HI}=\mathbb{E}_{P_{Z}}[\rho^{\rm HI}(\lambda,Z)]$ the average energy efficiency of high-priority users, and as $(\alpha_{1},\alpha_{2})$ the corresponding requirements. A hyperparameter is reliable if it meets both the queue occupancy and energy efficiency requirements for high-priority users, i.e., $\Lambda^{\rm rel}=\{\lambda\in\Lambda:Q^{\rm HI}(\lambda)\leq\alpha_{1}\cap\rho^{\rm HI}(\lambda)\geq\alpha_{2}\}$. An e-process for the hypothesis $\mathcal{H}_{i}:Q^{\rm HI}(\lambda_{i})>\alpha_{1}\cap\rho^{\rm HI}(\lambda_{i})<\alpha_{2}$ can be obtained as the minimum between an e-process for hypothesis $\mathcal{H}^{1}_{i}:Q^{\rm HI}(\lambda_{i})>\alpha_{1}$, and an e-process for the hypothesis $\mathcal{H}^{2}_{i}:\rho^{\rm HI}(\lambda_{i})<\alpha_{2}$.
@@ 6 Conclusion
We introduced aLTT, a novel framework for hyperparameter selection that implements data-dependent sequential testing via early termination. Unlike the existing LTT, which builds on p-value multiple hypothesis testing (MHT), aLTT is based on sequential MHT via e-processes (Xu et al., 2021). In practical scenarios, this results in more efficient and flexible calibration procedures that maintain statistical validity and the same discovery power as LTT by using only a fraction of testing rounds.
Potential extensions of the aLTT framework include the study of scenarios characterized by distribution shift, simulation-aided calibration, and the application of aLTT to the calibration of large language models.
@@ 7 Acknowledgments
This work was supported by the European Union’s Horizon Europe project CENTRIC (101096379). The work of Osvaldo Simeone was also supported by the Open Fellowships of the EPSRC (EP/W024101/1) and by the EPSRC project (EP/X011852/1).
@@ A.2.1 Benjamini-Hochberg Procedure
By (18), BH applies a larger threshold to hyperparameters $\lambda\in\Lambda$ that have larger p-values and are thus less likely to be reliable. If the p-values in set $\mathcal{P}$ are independent, the Benjamini-Hochberg (BH) procedure $\mathcal{A}^{\rm FDR}_{\rm BH}(\mathcal{P})$ guarantees FDR control at a level $\delta$ (Benjamini and Hochberg, 1995).
@@ A.2.2 Benjamini-Yekutieli Procedure
The Benjamini-Yekutieli (BY) procedure $\mathcal{A}^{\rm FDR}_{\rm BY}(\mathcal{P})$ is $\delta$-FDR-controlling (Benjamini and Yekutieli, 2001).
@@ A.2.3 E-Benjamini-Hochberg Procedure
Following the same basic principles underlying BH in (18), eBH applies a larger threshold to hyperparameters $\lambda\in\Lambda$ that have smaller e-values and are thus less likely to be reliable. The eBH procedure returns an $\delta$-FDR-controlling set even for arbitrarily dependent e-values $E^{t}_{(k)}$ without the need to adjust the error level $\delta$ as in the case of the p-value-based BH procedure (Xu et al., 2021).
@@ B.1.1 Effect of the Betting Strategy
In Section 5.1, we studied the TPR of aLTT under the approximate growth rate adaptive to the particular alternative (aGRAPA) betting strategy. In the following, we compare its performance against non-adaptive betting strategies, specifically the maximum bet (MaxBet) strategy, which sets the bet $\mu^{t}_{i}$ to the maximum allowed value of $1/\alpha$, and the unitary bet (UnitBet) strategy, where $\mu^{t}_{i}=1$. Additionally, we evaluate two alternative adaptive online betting strategies based on approximate wealth maximization (Waudby-Smith and Ramdas, 2024): lower-bound on the wealth (LBOW) and online Newton step (ONS).
@@ B.2 Reliable Hyperparameter Selection for Wireless Resource Allocation
B.2.1 Minimizing Average Delay under an Energy Efficiency Requirement
Here we consider the wireless resource allocation problem of Section 5.2 under a reliability requirement on the average energy efficiency. Specifically, we consider the population risk $R(\lambda)=\mathbb{E}_{P_{Z}}[\rho(\lambda,Z)]$, so that the set of reliable hyperparameters is given by $\Lambda^{\rm rel}=\{\lambda\in\Lambda:R(\lambda)\geq\alpha\}$ for the given target energy efficiency $\alpha$. After $T=1000$ testing steps, with each step consisting of a resource allocation episode of $K=2500$ frames, the hyperparameter selection scheme selects from the estimated subset of reliable hyperparameters $\hat{\Lambda}^{\rm rel}$, the hyperparameter $\hat{\lambda}$ that minimizes the empirical communication delay $\hat{T}_{\rm tx}(\lambda,\mathcal{D}^{T})=1/|\mathcal{D}^{T}|\sum_{Z\in\mathcal{D}^{T}}\hat{T}_{\rm tx}(\lambda,Z)$ based on all the data $\mathcal{D}^{T}$ collected across the testing steps.

### v2

@@ 5.2.1 Problem Definition
Figure 4: True positive rate of LTT and aLTT with $\epsilon$-greedy acquisition policy for $\epsilon\in\{0.25,0.5,0.75,0.95\}$ and non-adaptive acquisition. On the left panel, the prediction sets satisfy FWER control, while on the right panel they meet FDR requirements. In both cases, the tolerance level is $\delta=0.1$. Results are averaged over the tasks in (Honovich et al., 2022) that yield non-empty reliable prompts set.
Consider a target LLM $f(\cdot)$ that, when fed the concatenation $[\lambda,X]$ of an instruction $\lambda$ and data $X$, produces an output text $f([\lambda,X])$. For example, in the task of movie recommendation illustrated in Figure 1, each prompt $\lambda$ corresponds to an instruction to generate movie titles similar to those provided in the input list $X$; while the output $f([\lambda,X])$ represents a list of recommended movies. Following (Zhou et al., 2023; Zhang et al., 2023), the candidate set $\Lambda$ of instructions, corresponding to the hyperparameters to be tested, is generated using a separate LLM. In particular, we use the Llama3 8B Instruct LLM (Dubey et al., 2024) as the target model $f(\cdot)$, while the initial instruction set $\Lambda$ is generated through the Llama3.3 70B Instruct model (meta, 2025) using the forward generation mode detailed in (Zhou et al., 2023).
For each instruction $\lambda$, the ability of the prompted model $f([\lambda,X])$ to generate high-quality outputs for a test datum $Z=(X,Y)$ is measured by a task-dependent loss function $R(\lambda,Z)=\ell(f([\lambda,X]),Y)\in[0,1]$. Focusing on tasks from the instruction induction data set (Honovich et al., 2022), we specifically adopt the 0-1 loss tailored to the given task (Zhou et al., 2023). The goal is to determine a subset $\Lambda^{\rm rel}\subseteq\Lambda$ that complies with the condition (2) with a target execution error $\alpha=0.2$, and FWER and FDR requirements given by the tolerance parameter $\delta=0.1$.
@@ 5.2.2 Results
For the outlined prompt engineering problem, we evaluate aLTT using an $\epsilon$-greedy acquisition strategy with $\epsilon\in\{0.25,0.5,0.75,0.95\}$, as well as a non-adaptive acquisition strategy whereby the instruction to be tested is selected randomly and independently from prior testing rounds. All schemes employ an online Newton step betting strategy (Waudby-Smith & Ramdas, 2024) for the e-value (12).
Figure 5: Length of the shortest instruction in the predicted set of reliable hyperparameter $\hat{\Lambda}^{\rm rel}$ returned by LTT and aLTT. Instructions are tested under different accuracy requirements and with a fixed testing budget $T=2000$.
In Figure 4, we report the TPR of LTT and aLTT as a function of the number of rounds $t$. Results are averaged over tasks that yield a final non-empty set $\hat{\Lambda}^{\rm rel}$. Plots for individual tasks are included in Appendix B.3. Under both FDR and FWER control, aLTT reduces the number of testing rounds, or equivalently the number of LLM executions, required to achieve a given TPR. In particular, LTT produces a non-empty set only at the end of the testing process, attaining a final TPR less than half of the TPR achieved by aLTT with $\epsilon$-greedy acquisition strategy and $\epsilon=0.25$. Moreover, aLTT with $\epsilon=0.25$ identifies $50\%$ of the reliable instructions within the first $1000$ testing rounds, while the non-adaptive testing procedure identifies less than $10\%$ of the reliable instructions with the same number of rounds.
The improved efficiency of aLTT translates into superior performance in downstream tasks. To illustrate this point, we follow the described hyperparameter selection process with a post-selection phase that identifies a single hyperparameter $\hat{\lambda}$ from the estimated set $\hat{\Lambda}^{\rm rel}$ of reliable instructions (see Figure 1). Specifically, we adopt the shortest instruction $\lambda\in\hat{\Lambda}^{\rm rel}$. This post-selection criterion is motivated by the computational benefits of processing shorter prompts.
In Figure 5, we report the length of the selected instruction $\hat{\lambda}$ as a function of the target execution accuracy $1-\alpha$. Results are averaged over tasks with a final non-empty set $\hat{\Lambda}^{\rm rel}$, while plots for individual tasks are included in Appendix B.3. Stricter accuracy requirements are seen to reduce the number of discovered reliable instructions, leading to an increase in the minimal instruction length within $\hat{\Lambda}^{\rm rel}$. However, across all reliability levels $\alpha$, aLTT consistently delivers the shortest instructions, outperforming alternative schemes. This advantage is to be attributed to the data-adaptive acquisition strategy implemented by aLTT, which enables the discovery of a larger number of reliable instructions.
@@ 6 Conclusion
We introduced aLTT, a novel framework for hyperparameter selection that implements data-dependent sequential testing via early termination. Unlike the existing LTT, which builds on p-value multiple hypothesis testing (MHT), aLTT is based on sequential MHT via e-processes (Xu et al., 2021). In practical scenarios, including the problem of prompt engineering, this results in more efficient and flexible calibration procedures that maintain statistical validity and the same discovery power as LTT by using only a fraction of testing rounds.
Potential extensions of the aLTT framework include the study of scenarios characterized by distribution shift and simulation-aided calibration.
@@ 7 Acknowledgments
This work was supported by the European Union’s Horizon Europe project CENTRIC (101096379). The work of Osvaldo Simeone was also supported by the Open Fellowships of the EPSRC (EP/W024101/1) and by the EPSRC project (EP/X011852/1). Thanks also to the Advanced Research and Invention Agency (ARIA) for supporting broader foundational work in this space.
@@ A.2.1 Benjamini-Hochberg Procedure
By (18), BH applies a larger threshold to hyperparameters $\lambda\in\Lambda$ that have larger p-values and are thus less likely to be reliable. If the p-values in set $\mathcal{P}$ are independent, the Benjamini-Hochberg (BH) procedure $\mathcal{A}^{\rm FDR}_{\rm BH}(\mathcal{P})$ guarantees FDR control at a level $\delta$ (Benjamini & Hochberg, 1995).
@@ A.2.2 Benjamini-Yekutieli Procedure
The BY procedure $\mathcal{A}^{\rm FDR}_{\rm BY}(\mathcal{P})$ is $\delta$-FDR-controlling (Benjamini & Yekutieli, 2001).
@@ A.2.3 E-Benjamini-Hochberg Procedure
Following the same basic principles underlying BH in (18), eBH applies a smaller threshold to hyperparameters $\lambda\in\Lambda$ that have smaller e-values and are thus less likely to be reliable. The eBH procedure returns an $\delta$-FDR-controlling set even for arbitrarily dependent e-values $E^{t}_{(k)}$ without the need to adjust the error level $\delta$ as in the case of the p-value-based BH procedure (Xu et al., 2021).
@@ B.1.1 Effect of the Betting Strategy
In Section 5.1, we studied the TPR of aLTT under the approximate growth rate adaptive to the particular alternative (aGRAPA) betting strategy. In the following, we compare its performance against non-adaptive betting strategies, specifically the maximum bet (MaxBet) strategy, which sets the bet $\mu^{t}_{i}$ to the maximum allowed value of $1/\alpha$, and the unitary bet (UnitBet) strategy, where $\mu^{t}_{i}=1$. Additionally, we evaluate two alternative adaptive online betting strategies based on approximate wealth maximization (Waudby-Smith & Ramdas, 2024): lower-bound on the wealth (LBOW) and online Newton step (ONS).
@@
B.1.2 Quantile Risk Control
Figure 7: True positive rate of LTT and aLTT with $\epsilon$-greedy acquisition policy for $\epsilon\in\{0.25,0.5,0.75,0.95\}$ and non-adaptive acquisition. On the left panel, the prediction sets satisfy FWER control while on the right FDR control. In both cases, a policy is considered reliable if the quantile risk at level $q=0.1$ is lower than $\alpha=0.57$, and the tolerance level is $\delta=0.1$.
In this experiment, we consider the same set-up of Section 5.1 with the caveat that, instead of identifying policies with a large enough average reward, we focus on policies with a small enough quantile risk — a risk definition that explicitly captures safety (Dabney et al., 2018; Farzaneh et al., 2024). The quantile risk of a hyperparameter $\lambda$ at a level $q$ is defined as
$R_{q}(\lambda)=\inf_{r}\{r:\Pr[R(\lambda,Z)\leq r]>1-q\}.$ (21)
Controlling the quantile risk ensures guarantees on the average fraction of times the policy risk exceeds the level $R_{q}(\lambda)$. For the RL setting, we deal with rewards, so the inequalities in the definition are reversed.
We aim to identify policies with a quantile risk at the level $q=0.1$ that is smaller than $\alpha=0.57$. In Figure 7, we report the true positive rate of aLTT using the $\epsilon$-greedy acquisition policy and LTT. aLTT significantly outperforms LTT, identifying up to 75% of the policies that satisfy the given reliability requirement within the maximum testing round, $T$. In contrast, LTT is unable to identify any policy due to its non-adaptive testing strategy.
@@ B.2 Reliable Hyperparameter Selection for Wireless Resource Allocation
B.2.1 Problem Definition
[…]
B.2.2 Results
We study a setting with differentiated service requirements. In particular, we enforce a constraint on the average queue occupancy and energy efficiency for high-priority users, while making a best-effort attempt at minimizing the energy-delay product for low-priority users. Writing as $Q^{\rm HI}(\lambda)=\mathbb{E}_{P_{Z}}[Q^{\rm HI}(\lambda,Z)]$ the average queue occupancy of high-priority users, and as $\rho^{\rm HI}=\mathbb{E}_{P_{Z}}[\rho^{\rm HI}(\lambda,Z)]$ the average energy efficiency of high-priority users, and as $(\alpha_{1},\alpha_{2})$ the corresponding requirements. A hyperparameter is reliable if it meets both the queue occupancy and energy efficiency requirements for high-priority users, i.e., $\Lambda^{\rm rel}=\{\lambda\in\Lambda:Q^{\rm HI}(\lambda)\leq\alpha_{1}\cap\rho^{\rm HI}(\lambda)\geq\alpha_{2}\}$. An e-process for the hypothesis $\mathcal{H}_{i}:Q^{\rm HI}(\lambda_{i})>\alpha_{1}\cap\rho^{\rm HI}(\lambda_{i})<\alpha_{2}$ can be obtained as the minimum between an e-process for hypothesis $\mathcal{H}^{1}_{i}:Q^{\rm HI}(\lambda_{i})>\alpha_{1}$, and an e-process for the hypothesis $\mathcal{H}^{2}_{i}:\rho^{\rm HI}(\lambda_{i})<\alpha_{2}$. Note that any e-merging function can be applied to ensure validity over the joint hypothesis (see, e.g., (Vovk & Wang, 2021; Ramdas & Wang, 2024)), we choose minimum for simplicity.
B.2.3 Minimizing Average Delay under an Energy Efficiency Requirement
Now we consider the wireless resource allocation problem under a reliability requirement on the average energy efficiency. Specifically, we consider the population risk $R(\lambda)=\mathbb{E}_{P_{Z}}[\rho(\lambda,Z)]$, so that the set of reliable hyperparameters is given by $\Lambda^{\rm rel}=\{\lambda\in\Lambda:R(\lambda)\geq\alpha\}$ for the given target energy efficiency $\alpha$. After $T=1000$ testing steps, with each step consisting of a resource allocation episode of $K=2500$ frames, the hyperparameter selection scheme selects from the estimated subset of reliable hyperparameters $\hat{\Lambda}^{\rm rel}$, the hyperparameter $\hat{\lambda}$ that minimizes the empirical communication delay $\hat{T}_{\rm tx}(\lambda,\mathcal{D}^{T})=1/|\mathcal{D}^{T}|\sum_{Z\in\mathcal{D}^{T}}\hat{T}_{\rm tx}(\lambda,Z)$ based on all the data $\mathcal{D}^{T}$ collected across the testing steps.
@@ B.2.3 Minimizing Average Delay under an Energy Efficiency Requirement
B.3 Reliable Prompt Selection for Natural Language Processing
In Figures 4 and 5, we report the performance averaged over ten tasks from the instruction induction dataset (Honovich et al., 2022). Complementing the results in the main text, here we provide a breakdown of the methods’ performance across the ten tasks. In Figure 11, we show the TPR of the schemes considered in Section 5.2.2 under FDR and FWER control with tolerance $\delta=0.1$. Across all tasks, aLTT consistently outperforms LTT, which returns empty sets for many tasks.
In Figure 12, we plot the length of the shortest instruction in the predicted set $\hat{\Lambda}^{\text{rel}}$ returned by the different hyperparameter selection schemes as a function of the execution accuracy $1-\alpha$. If a prediction scheme returns an empty predicted hyperparameter set, we set the instruction length to the length of the longest instruction in the candidate set $\Lambda$. For all tasks, aLTT can identify shorter instructions compared to LTT under the same testing budget.
Figure 11: Breakdown of the average performance of methods presented in Figure 4 across the 10 tasks in the instruction induction dataset (Honovich et al., 2022) for which a non-empty hyperparameter sets $\hat{\Lambda}^{\text{rel}}$ is returned. Across all tasks, aLTT delivers the highest true positive rate.
Figure 12: Length of the shortest instruction in $\hat{\Lambda}^{\text{rel}}$ as a function of target execution accuracy $1-\alpha$ for each task in the instruction induction dataset (Honovich et al., 2022) for which the methods presented in Figure 5 return non-empty hyperparameter sets $\hat{\Lambda}^{\text{rel}}$.